Category Archives: HPF

This is for the HET instrument Habitable-zone Planet Finder

The week of October 16

The exciting news this week has been the arrival of the Habitable Planet Finder (HPF). This is the first of our new high resolution instruments and an instrument well suited to working in bright moon conditions. The HPF was designed and built by our Penn State partners and arrived on the 16th.

This instrument is designed for extremely high precision spectroscopy capable of detecting the reflex motion of stars as small earth sized planets go around them. To achieve that precision the instrument is housed in our temperature controlled basement at the HET inside a temperature controlled room inside a large temperature controlled vacuum chamber. All of these efforts allow them to control the temperature of the optics of the instrument at a level of 0.001 degrees Celsius.

After very carefully cleaning the enclosure that will house the vacuum chambered instrument the HPF team

was able to open their instrument and after a very through inspection proudly announced that they have just as many pieces of glass as they did in the assembly lab at Penn State (an optics joke). After a few final checks and the inclusion of their single moving part inside of the spectrograph they sealed up the vacuum chamber which, if things continue to go very well, may remain sealed for several years to come. The process of pumping the vacuum out of the large chamber took the rest of the weekend.

In the coming days and weeks the HPF team will monitor its stability, install the laser metrology system and get the systems ready for on-sky commissioning.

Quite an exhausting and exciting week at the HET!

The week of October 2nd

This week we are pleased to announce that a new VIRUS unit was installed in side two of the VIRUS enclosure. This brings us to 22 VIRUS units or 44 spectrographs. We also took a little time in the last engineering run to add on some valves to the vacuum fittings which will allow us to cold pump on the VIRUS units which takes far less time to do than to warm up and then repump which was our older methodology. Keeping 22 VIRUS units going is starting to be a little easier but still takes a lot of management.

In addition to the work on VIRUS, we have also installed in the coherent fiber bundles for the HPF. These coherent fiber bundles will be used to setup stars on HPF science fibers. HPF will arrive in the coming weeks and we are very excited to get our first high resolution instrument on sky in the coming months.

The week of Dec 7

Two weeks ago was the HET Board of Director’s meeting in Penn State. The meeting lasted two days and the Board got status reports from HET operations and each of the instrument teams. The main news is that that LRS2 is expecting to reach first science in the first few months of 2016, HRS2 is going to start commissioning before Summer 2016 and VIRUS units are going to be coming in over the next 9 months. The Board was encouraged by the progress being made and hopes that we can continue the pace. They were also pleased to see the progress being made on HPF and were impressed with the clean room facility tour they were given. No major changes or action items were reported by the Board.

Once the commissioning team returned from Happy Valley, we went right back to work and were able to push along the closure of one of our major metrology loops, the guide probes. I am pleased to report that we are able to guide at any telescope Az for full trajectories with the probes at any position within their range.

The week of January 26

Happy New Year! The blogger is back from Winter break and travels to January Science meetings.

We will start with the bad news. The tests on the corrector have determined that there is significant asymmetries. These are asymmetries beyond what were found and reported to the board of directors in the December meeting. Those aberrations were going to be removed by changing the plate glass that was going in to seal the bottom of the corrector into a fifth optical element with some small power. That optic has been fabricated and installed in the corrector. The newly discovered astigmatism was found off axis and might have an impact on image quality at the edge of the field of view. To further investigate this optical issue we are adding a few months to the delivery date of the corrector. Instead of being delivered in late January the new delivery date is expected to be early May and significant on-sky commissioning during the rainy season perhaps pushing us into September.

On a more positive note the 2nd VIRUS enclosure has arrived and installed.

Panoramic views showing the VIRUS Enclosures being installed.  The top shows a view from in-front of the mirror and the bottom shows a view from behind the mirror.  Both are taken from the catwalk.

Panoramic views showing the VIRUS Enclosures being installed. The top shows a view from in-front of the mirror and the bottom shows a view from behind the mirror. Both are taken from the catwalk.

There is a lot of plumbing and electrical work required to finish the installation. This is our highest priority.

During the long break the Remote Thermal Area contractors have installed the glycol chiller and are in the process of commissioning it with glycol. This nearly completes the Remote Thermal Area project. Only a few punch list items remain.

The HPF doors also arrived and have been installed. The HPF is now thermally isolated from the spectrograph room. The Penn State Team will monitor the temperatures inside and outside of their enclosure to see if our new Mitsubishi units can hold the temperature to a tight enough tolerance for their specifications. Meanwhile the insulating panels for the HRS II have arrived and will be installed in the coming month. These panels along with an active feed back heater inside the enclosure should allow the HRS II to be held to extremely tight tolerances, < 0.1 C.

The mirror team has started mirror swaps and started using the Strip and Wash room to remove old mirror coatings and prep them for the coating chamber. Our goal is to get the team up to 4 mirrors swaps every 3 weeks.

The week of Nov 10

The biggest change the occurred this week was the completion of the HPF enclosure. A team came out from Penn State to lead the efforts and after two days the enclosure was completed and the last day was spent wiring up temperature sensors to monitor how the systems reacts to seasonal changes in the HET basement.

Newly installed HPF enclosure next to the HRS enclosure.

Newly installed HPF enclosure next to the HRS enclosure.

Our electrical team completed hook up of power cables to distribution boxes in Virus annex this week. They hoisted and installed the Virus equipment rack in the annex as well. We hope to have all of the electrical work done for the first enclosure by the end of next week when a programmer comes from Austin work work on the PLCs.

This week we found a problem with loose screws on lower X. We shut the tracker down for the week and an engineer came out from Austin to inspect the system for any damage. He found that many of the screws were loose and some of the shims were missing. This was likely caused by a fairly strong vibration incident that we had when we were commissioning the TCS timing loops. The engineers and mechanical team got the X drive shimmed and aligned again and all of the crews have been tightened down with torque wrenches. While the engineer was out here he worked on the rho stage which was hitting the proximity sensors early due to the non-flatness of the switch plate. They re-installed it with shims to adjust flatness and installed RHO limit switches and set the clearance.

The initial commissioning of the Strip and Wash Room began this week. A mirror segment was put through the new ultra-pure water cleaning. It is not entirely surprising that a number of small leaks were found around the mirror and passage way doors. The rest of the week was spent finding leaks and diverting water drainage paths. Our electrical engineer has nearly completed the PID controller and pressure display that will show when the room is in positive pressure and negative pressure.

As an interesting diversion this week we had the Director on site with several photographers taking pictures of him and the facility. Some of the shots were done from a helicopter while others were done from various catwalks or man-lifts. We all look forward to seeing these glamour shots.

The week of Nov 3

The in-line fans for the K-hut and Strip and Wash room are now in place and working. This means that we can begin the process of commissioning the strip and Wash room. At the moment we still need a few more safety indicators on if the room has negative or positive pressure and to get the hot and cold water nozzles in place. This work seems to be moving forward at a good pace.

The replacement of the old truss sensors is moving forward. As mentioned before the old sensors have been removed and the new ones are being assembled.

Lift 3rd VSS platform into place, test fit, secured in place in prep for install.
Moved upper VSS platform into bay area, painted modified area and moved onto dome floor.
Work around the VIRUS Support Structure (VSS) continues. We have made some modifications to the work platforms and three of those have been test fit, secured in place and in preparation for the installation in two weeks. This week we did some work on possible coupling between the VSS and the structure. Our concern was that wind shake on the VSS would cause the mirrors to move. We did a number of test at the CCAS tower involving people physically moving the currently empty VSS enclosures. The concern seems to be warranted as the FWHM of the induced movement is about a quarter of an arc-second. We will look into identifying the members that cause the most coupling between the two and mitigate the motion.

The software team out here has been working on the DMI and Tip/Tilt to get it realigned with the new mount models that were installed last week.

The big news this week is the arrival of the HPF enclosure. Despite some very thick fog and misty rain we were able to get the truck unloaded and placed in the spectrograph room without tearing up the newly prepared floor. The PSU assembly team will arrive late next week.

The week of Oct 27

The contractors continue to work through the Remote Thermal Area punch list including
repairing sheet rock, cleaning up building penetrations, clean up material and reinstalling the roof gutter down spouts. We also have contractors out working on the balance of the in-line fans for the strip and wash duct and the K-hut duct.

We received another shipment of VIRUS Platforms/Ladders;
it contained the platform side supports for both VIRUS right and VIRUS left and a top work platform for one of the VIRUS enclosures. We have been doing test fits of these enclosures in preparation for installing all of them in two weeks.

John Good, mechanical engineer from Austin, was here for the last two weeks to work with the laser tracker. The first days were spent confirming the current mount model then making refinements to it. We now believe we have the best mount model that can be obtained with this laser tracker.

Some work at night was done this week. We went on sky to test the mount model, see if the position of the Celestron Alignment Telescope (CAT) was good enough for future testing. We found the RA, DEC and rho offsets are now working and that the center of rotation falls at the corner of the CAT. Tests of the geosynchronous satellite tracking software within TCS suggests that it is probably working although the CAT is just to small to see these 11-13th mag objects. We believe we are sufficiently ready for any future metrology tests. With this alignment out of the way we put the CCAS tower mirror alignment system back together and have confirmed that all of the components are well aligned and working with alignment to better than 0.5″.

Work in the spectrograph room continues. One of our TOs finished the touch-ups required after the contractors scratched the paint in the area where the HPF enclosure will go. With that done we are ready for the HPF enclosure to arrive next week. We also received the mounts for the HRS sensors. These will be part of the system that monitors and controls the HRS temperature to improve performance in velocity stability.

The really big news is that we installed the large liquid nitrogen tank behind the K-hut. This required days of preparation and coordination between the HET, Crane services, Praxair and Midwest Cryo teams. The two cranes made quick work of putting the crane in place and by lunch the tank was bolted down. In the two following days nearly all of the plumbing to the K-hut vaporizer and the filling location were installed.

DSC_0061

The week of March 31st

In the past week the HPF team from Penn State came out to remove the last components of the MRS and install the calibration enclosure for the HPF. This work went well and their enclosure is now in place. They have approved the floor paint which will go below the HPF main enclosure when they bring it out in late 2014 or early 2015.

We also made a jump forward with the mirror Strip And Wash (SAW) room. This room is connected to the mirror coating chamber and now we can do a final ZrO strip of the surface and a ultra pure water rinse and take it directly into the coating room without leaving a clean room area. This SAW room is now being used in a very manual mode and we hope to automate some components in the future.

This past week we made some small improvements in the tracker performance but we continue to have problems with the encoder. We will continue to work on this problem this week.

This week may be a little slower until we can get the problems with the tracker fixed. This will give the operations support team a chance to catch up with maintenance issues and make progress on other projects like the mirror washing room.

We also may have some contractors for Veliz Construction on site to begin the process of aluminum taping of the control building roof. This will reduce its emissivity and should help it from super-cooling at night and sending cold air into the dome. Unfortunately for Veliz the wind does not look like it will give them much of a break to actually get the work done.

The week of March 24

It has been slightly slow going with the new track electronics. We have found that some communications are flaky (bad mil-spec connections) and many of the motors require significant retuning. The PFIP test mass is in place but we want to make sure we have all motions largely tuned and reliable before we pull down the test mass and replace it with the hexapod assembly and the PFIP frame.

We have had several pieces of bad new from the University of Arizona group who are assembling the HET corrector. At the moment we have had a slip in the delivery date of three weeks and may see further slips in the near future depending upon the nature of the problems they have encountered.

Later this week the HPF team from Penn State will arrive and will be shipping out the last of the MRS and FIF components and assembling the calibration enclosure for the HPF.